# Time-varying Vector Field Compression with Preserved Critical Point Trajectories

### v2

@@ VII-E Encoding Efficiency
In our pipeline, two factors dominate the final bit rate: the fixed error bound after quantization $\{Q_{\xi}\}$that guarantees critical-point tracking after quantization, and the distribution of the quantization indices $\{Q_{d}\}$. The error bound is analytically derived from the requirement and is therefore not a tunable lever for further compression. Consequently, our effort focuses on shaping the index distribution via better prediction so that the subsequent entropy and lossless coders operate more efficiently. The rate gains of MoP come from two layers: (1) Zero-order statistics. Let $p(k)$ denote the PMF of the folded quantization index (central bins correspond to $|q|\leq 1$ under our toward-zero mapping). The ideal per-sample cost of Huffman coding is $H_{0}(p)=-\sum_{k}p(k)\log_{2}p(k)$. By concentrating probability mass near the center, MoP reduces $H_{0}(p)$ compared with 3DL and the SL predictor. In Figure 9, the MoP method exhibits a markedly sharper central peak than 3DL and SL. Meanwhile, the tail CCDF illustrates the relative proportions of data in both tails. It can be seen that MoP exhibits a smaller proportion beyond the range [3,3][-3,3], suggesting a higher concentration of data around the center which directly translates into shorter Huffman codewords and a lower first-order rate. (2) Higher-order structure. After entropy coding, the bytestream is fed to a backend lossless compressor (Zstd), which exploits repetition and context (LZ77-style matching) beyond zero-order statistics. Figure 10 shows the complementary cumulative distribution function(CCDF) of the run lengths for the center symbol obtained from three different predictors: 3DL, SL, and MoP. The x-axis represents the run length $L$, while the y-axis denotes the probability $P(run\geq L)$, i.e., the proportion of continuous run length is greater than or equal to $L$. This figure provides a statistical view of how long the correctly predicted regions persist continuously across space and time. In other words, a fatter tail corresponds to more regular and persistent higher-order structures. According to the figure, we can observe that both SL and MoP are able to capture more longer run-lengths compared with 3DL. In addition, the upper-right panel of the figure shows the corresponding means and multiple percentiles for the three methods, which further confirm this observation. Since zstd leverages repeated substrings, a longer zero-run yields longer matches and thus additional savings on top of Huffman, further improving the overall compression ratio.
@@
Figure 12: Visualization of critical point trajectories in the FS dataset.
@@
VII-F Comparison with State of the Art
@@
We present the quantitative results for all datasets in Table II. These results include an ablation study of the predictor component, in which we evaluate three configurations of our method. The Default configuration uses the standard 3D Lorenzo predictor, the SL configuration replaces it with our proposed Semi-Lagrangian predictor, and the MoP configuration adaptively selects between the two predictors. All reported experimental runtimes correspond to serial execution without any parallelization. To ensure a fair comparison, we adjusted the parameters of SZ3 and cpSZ so that their compression ratios are as close as possible to that of our method in the default (3DL) configuration. According to the results, since SZ3 does not provide any guarantee of preserving critical points, it consequently fails to ensure the correctness of critical point trajectories. Although cpSZ can preserve the critical points at each individual time step, it cannot guarantee the consistency of critical points within a time slab, which leads to inconsistent trajectory counts and the emergence of false cases within the slab. Among all the lossless compression frameworks, FPZIP achieves the best performance, with a maximum compression ratio of $4.06$. The results of GZIP and ZSTD are comparable, with compression ratios ranging between $1.1$ and $1.3$.
@@
On the other hand, all three of our proposed methods effectively preserve critical point trajectories. Among them, MoP consistently outperforms 3DL across all four datasets, with particularly notable improvements on the CFVKV and FS datasets, achieving $72.29\times$ and $124.48\times$ compression ratios, respectively. Furthermore, when using SL alone as the predictor, it surpasses 3DL on three of the datasets with a maximum improvement of 32%, and performs comparably to 3DL on the SCF dataset. Overall, among all compressors capable of preserving critical point trajectories, our method achieve up to $56.07\times$ in compression ratio compared to the state-of-the-art compressors across all datasets. We observe that in the SCF dataset, a large fraction of vertices have $d_{\infty}$ significantly exceeding $d_{\max}$. Specifically, 85.4% and 69.0% of the vertices triggered the adaptive substepping mechanism in SL and MoP methods, respectively, with an average of 4.1 and 3.7 substeps. no significant increase in runtime was observed,, indicating that the adaptive substepping mechanism does not introduce significant overhead in practice.
@@
Figure. 11 shows the compression and decompression time across all the compressors. ZSTD is the fastest among all lossless compressors, while ZFP achieves the highest speed among lossy compressors. Compared with cpSZ, our methods are more time-consuming because they aim to preserve trajectories, which requires examining up to 36 surrounding triangles for each point, including pre-computation of critical points and derivation of error bounds. Among our three proposed methods, 3DL and SL exhibit similar compression times, whereas MoP requires more time due to the additional block partitioning, sampling, and micro-encoding processes used for score calculation. For decompression time, 3DL and SL perform almost identically, while MoP incurs a slightly higher cost because it must switch predictors across different data blocks.
@@
Although our algorithm is slower than some existing lossy compressors, it uniquely preserves the critical point trajectories throughout the compression–decompression process, ensuring topological consistency in the reconstructed vector fields. Meanwhile, the framework is highly amenable to parallelization, which can substantially improve its performance. Moreover, our SL predictor achieves an even higher degree of parallelism than the traditional Lorenzo predictor, since it is embarrassingly parallel across all data within frame $t$ without depending on any current-frame neighbors. Consequently, existing studies [cusz, cuszp, tian2021cuszplus, liu_tian_wu2024cuszi] have demonstrated that prediction-based compressors can further accelerate throughput significantly on GPU. In addition, the decompression process in our framework is relatively fast, and since data are typically compressed once but decompressed multiple times, the additional compression overhead has little impact on practical usability.
@@ VII-G Memory Consumption
We further observe that FPZIP, ZFP, and SZ3 require approximately $1\times$, $3\times$, and $4\times$ the data size in memory, respectively, while both cpSZ-SoS and our method require slightly more than $5\times$. The additional memory usage of our method mainly stems from two factors: the use of int64 representations to support the SoS mechanism, which adds one extra data-size overhead, and an auxiliary hash map for maintaining critical point information. We note that SoS is not inherently tied to this implementation choice and could alternatively be realized using arbitrary-precision libraries (e.g., GMP) with reduced storage overhead at the cost of higher computational expense. Our current design favors computational efficiency and simplicity in implementation.
@@ VII-H Qualitative Visualizations
Critical point trajectories are extracted using the algorithm provided in FTK [ftk]. We overlay the line integral convolution (LIC) texture and the corresponding velocity field across time to display the critical point trajectories. As shown in the visualizations, our method completely preserves the original critical point trajectories. Both cpSZ and SZ3 distort certain trajectories in various ways, including but not limited to: (1) discontinuous trajectories accompanied by small oscillations, exhibiting zigzag patterns in the trajectories, which indicate slight positional deviations of critical points in the decompressed data; and (2) short-lived artifacts, manifested as ”small bubbles” that suddenly appear, split and merge within a short time slab, where these critical points do not exist in the original data.
@@ VIII Conclusion and Future Works
[…]
AI-Generated Content Acknowledgement

### v3

@@ VII-E Encoding Efficiency
In our pipeline, two factors dominate the final bit rate: the fixed error bounds after quantization, $\{Q_{\xi}\}$, which guarantee critical-point tracking, and the distribution of the quantization indices, $\{Q_{d}\}$. The error bound is analytically derived from the requirement and is therefore not a tunable lever for further compression. Consequently, our effort focuses on shaping the index distribution via better prediction so that the subsequent entropy and lossless coders operate more efficiently. The rate gains of MoP come from two layers: (1) Zero-order statistics. Let $p(k)$ denote the PMF of the folded quantization index (central bins correspond to $|q|\leq 1$ under our toward-zero mapping). The ideal per-sample cost of Huffman coding is $H_{0}(p)=-\sum_{k}p(k)\log_{2}p(k)$. By concentrating probability mass near the center, MoP reduces $H_{0}(p)$ compared with 3DL and the SL predictor. In Figure 9, the MoP method exhibits a markedly sharper central peak than 3DL and SL. Meanwhile, the tail CCDF illustrates the relative proportions of data in both tails. It can be seen that MoP exhibits a smaller proportion beyond the range [3,3][-3,3], suggesting a higher concentration of data around the center which directly translates into shorter Huffman codewords and a lower first-order rate. (2) Higher-order structure. After entropy coding, the bytestream is fed to a backend lossless compressor (ZSTD), which exploits repetition and context (LZ77-style matching) beyond zero-order statistics. Figure 10 shows the complementary cumulative distribution function (CCDF) of the run lengths for the center symbol obtained from three different predictors: 3DL, SL, and MoP. The x-axis represents the run length $L$, while the y-axis denotes the probability $P(run\geq L)$, i.e., the proportion of continuous run length is greater than or equal to $L$. This figure provides a statistical view of how long the correctly predicted regions persist continuously across space and time. In other words, a fatter tail corresponds to more regular and persistent higher-order structures. According to the figure, we can observe that both SL and MoP are able to capture more longer run-lengths compared with 3DL. In addition, the upper-right panel of the figure shows the corresponding means and multiple percentiles for the three methods, which further confirm this observation. Since ZSTD leverages repeated substrings, a longer zero-run yields longer matches and thus additional savings on top of Huffman, further improving the overall compression ratio.
@@
Figure 12: Visualization of critical-point trajectories in the FS dataset.
@@
VII-F Comparison with State-of-the-Art Methods
@@
We present the quantitative results for all datasets in Table II. The results include an ablation study of the predictor component, in which we evaluate three configurations of our method. The default configuration employs the standard 3D Lorenzo predictor (3DL); the SL configuration replaces it with the proposed semi-Lagrangian predictor; and the MoP configuration adaptively selects between the two predictors. All reported experimental runtimes correspond to serial execution without any parallelization. To ensure a fair comparison, we adjusted the parameters of SZ3 and cpSZ so that their compression ratios are as close as possible to that of our method in the default (3DL) configuration. According to the results, since SZ3 does not provide any guarantee of preserving critical points, it consequently fails to ensure the correctness of critical-point trajectories. Although cpSZ can preserve the critical points at each individual time step, it cannot guarantee the consistency of critical points within a time slab, which leads to inconsistent trajectory counts and the emergence of false cases within the slab. Among all the lossless compression frameworks, FPZIP achieves the best performance, with a maximum compression ratio of $4.06$. The results of GZIP and ZSTD are comparable, with compression ratios ranging between $1.1$ and $1.3$.
@@
On the other hand, all three of our proposed methods effectively preserve critical-point trajectories. Among them, MoP consistently outperforms 3DL across all four datasets, with particularly notable improvements on the CFVKV and FS datasets, achieving $72.29\times$ and $124.48\times$ compression ratios, respectively. Furthermore, when using SL alone as the predictor, it surpasses 3DL on three of the datasets with a maximum improvement of 32%, and performs comparably to 3DL on the SCF dataset. Overall, among all compressors capable of preserving critical-point trajectories, our method achieve up to $56.07\times$ in compression ratio compared to the state-of-the-art compressors across all datasets. We observe that in the SCF dataset, a large fraction of vertices have $d_{\infty}$ significantly exceeding $d_{\max}$. Specifically, 85.4% and 69.0% of the vertices trigger the adaptive substepping mechanism in SL and MoP methods, respectively, with an average of 4.1 and 3.7 substeps. No significant increase in runtime has been observed, indicating that the adaptive substepping mechanism does not introduce significant overhead in practice.
@@
Figure 11 shows the compression and decompression time across all the compressors. ZSTD is the fastest among all lossless compressors, while ZFP achieves the highest speed among lossy compressors. Compared with cpSZ, our methods are more time-consuming because they aim to preserve trajectories, which requires examining up to 36 surrounding triangles for each point, including pre-computation of critical points and derivation of error bounds. Among our three proposed methods, 3DL and SL exhibit similar compression times, whereas MoP requires more time due to the additional block partitioning, sampling, and micro-encoding processes used for score calculation. For decompression time, 3DL and SL perform almost identically, while MoP incurs a slightly higher cost because it must switch predictors across different data blocks.
@@
Although our algorithm is slower than some existing lossy compressors, it uniquely preserves the critical-point trajectories throughout the compression–decompression process, ensuring topological consistency in the reconstructed vector fields. Meanwhile, the framework is highly amenable to parallelization, which can substantially improve its performance. Moreover, our SL predictor achieves an even higher degree of parallelism than the traditional Lorenzo predictor, since it is embarrassingly parallel across all data within frame $t$ without depending on any current-frame neighbors. Consequently, existing studies [cusz, cuszp, tian2021cuszplus, liu_tian_wu2024cuszi, songceresz, qiu2025] have demonstrated that prediction-based compressors can further accelerate throughput significantly on GPUs and other hardware accelerators. In addition, the decompression process in our framework is relatively fast, and since data are typically compressed once but decompressed multiple times, the additional compression overhead has little impact on practical usability.
@@ VII-G Memory Consumption
We further observe that FPZIP, ZFP, and SZ3 require approximately $1\times$, $3\times$, and $4\times$ the data size in memory, respectively, while both cpSZ-SoS and our method require slightly more than $5\times$. The additional memory usage of our method mainly stems from two factors: the use of int64 representations to support the SoS mechanism, which adds one extra data-size overhead, and an auxiliary hash map for maintaining critical-point information. We note that SoS is not inherently tied to this implementation choice and could alternatively be realized using arbitrary-precision libraries (e.g., GMP) with reduced storage overhead at the cost of higher computational expense. Our current design favors computational efficiency and simplicity in implementation.
@@ VII-H Qualitative Visualizations
Critical-point trajectories are extracted using the algorithm provided in FTK [ftk]. We overlay the line integral convolution (LIC) texture and the corresponding velocity field across time to display the critical-point trajectories. As shown in the visualizations, our method completely preserves the original critical-point trajectories. Both cpSZ and SZ3 distort certain trajectories in various ways, including but not limited to: (1) discontinuous trajectories accompanied by small oscillations, exhibiting zigzag patterns in the trajectories, which indicate slight positional deviations of critical points in the decompressed data; and (2) short-lived artifacts, manifested as “small bubbles” that suddenly appear, split and merge within a short time slab, where these critical points do not exist in the original data.
@@ VIII Conclusion and Future Works
In this paper, we address the problem of preserving critical-point trajectories during the lossy compression of 2D time-varying vector field data. We introduce a semi-Lagrangian–based predictor that achieves up to $4.31\times$ higher compression ratios than the Lorenzo predictor on advection-dominated datasets. Furthermore, we design a Mixture-of-Predictors module that adaptively selects the most suitable predictor for each local data block based on its spatial characteristics. This adaptive strategy effectively reduces residual entropy while maintaining high compression performance, achieving up to $56.07\times$ higher compression ratios than state-of-the-art lossless compressors in our experiments. In future work, we plan to further improve the efficiency of the framework, and to develop a GPU-accelerated implementation to achieve higher performance.
Acknowledgment
This work was partially supported by grants from NSF OAC-2311756, OAC-2442627, OAC-2504255, OAC-2313122, OAC-2313123, OAC-2313124, OIA-2327266, and OAC-2311878. We thank the University of Kentucky Center for Computational Sciences and Information Technology Services Research Computing for its support and use of the Lipscomb Compute Cluster, Morgan Compute Cluster, and associated research computing resources.
AI-Generated Content Acknowledgment